The value of a new car decreases by 20% in the first year and by 15% in the second year.Calculate the total percentage decrease in its value in the next two years.
Answer by ikleyn(52943)

.
Multiplying by 0.8 = 1-0.2 is equivalent decreasing by 20%.

So, new value of the car in the first year is 0.8*P, where P is the original price.



Multiplying by 0.85 = 1-0.15 is equivalent decreasing by 15%.

So, the value of the car in the two years is 0.8*0.85*P = 0.68*P.


Multiplying by 0.68 is equivalent decreasing the price by 1-0.68 = 0.32 = 32%.